# CellFlow: Simulating Cellular Morphology Changes via Flow Matching

### v1

@@ Abstract
Building a virtual cell capable of accurately simulating cellular behaviors in silico has long been a dream in computational biology. We introduce CellFlow, an image-generative model that simulates cellular morphology changes induced by chemical and genetic perturbations using flow matching. Unlike prior methods, CellFlow models distribution-wise transformations from unperturbed to perturbed cell states, effectively distinguishing actual perturbation effects from experimental artifacts such as batch effects—a major challenge in biological data. Evaluated on chemical (BBBC021), genetic (RxRx1), and combined perturbation (JUMP) datasets, CellFlow generates biologically meaningful cell images that faithfully capture perturbation-specific morphological changes, achieving a 35% improvement in FID scores and a 12% increase in mode-of-action prediction accuracy over existing methods. Additionally, CellFlow enables continuous interpolation between cellular states, providing a potential tool for studying perturbation dynamics. These capabilities mark a significant step toward realizing virtual cell modeling for biomedical research.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Overview of CellFlow. (a) Objective. CellFlow aims to predict changes in cell morphology induced by chemical or gene perturbations in silico. In this example, the perturbation effect reduces the nuclear size. (b) Data. The dataset includes images from high-content screening experiments, where chemical or genetic perturbations are applied to target wells, alongside control wells without perturbations. Control wells provide prior information to contrast with target images, enabling the identification of true perturbation effects (e.g., reduced nucleus size) while calibrating non-perturbation artifacts such as batch effects—systematic biases unrelated to the perturbation (e.g., variations in color intensity). (c) Problem formulation. We formulate the task as a distribution-to-distribution problem (many-to-many mapping), where the source distribution consists of control images, and the target distribution contains perturbed images within the same batch. (d) Flow matching. CellFlow employs flow matching, a state-of-the-art generative approach for distribution-to-distribution problems. It learns a neural network to approximate a velocity field, continuously transforming the source distribution into the target by solving an ordinary differential equation (ODE). (e) Results. CellFlow significantly outperforms baselines in image generation quality, achieving lower Fréchet Inception Distance (FID) and higher classification accuracy for mode-of-action (MoA) predictions.
@@
In this work, we introduce CellFlow, an image-generative model that simulates how cellular morphology changes in response to chemical or genetic perturbations (Figure 1a). CellFlow’s key innovation is formulating cellular morphology prediction as a distribution-to-distribution learning problem, and leveraging flow matching (Lipman et al., 2023), a state-of-the-art generative modeling technique designed for distribution-wise transformation, to solve this problem.
@@
Specifically, cell morphology data are collected through high-content microscopy screening, where images of control and perturbed cells are captured from experimental wells across different batches (Figure 1b). Control wells, which receive no drug treatment or genetic modifications, play a crucial role in providing prior information and serving as a reference to distinguish true perturbation effects from other sources of variation. They help calibrate non-perturbation factors, such as batch effects—systematic biases unrelated to perturbations, including variations in color or intensity, akin to distribution shifts in machine learning. Properly incorporating control wells is essential for capturing actual perturbation effects rather than artifacts, yet many existing methods overlook this aspect (Yang et al., 2021; Navidi et al., 2024; Cook et al., 2024). To address this, we frame cellular morphology prediction as a distribution-to-distribution mapping problem (Figure 1c), where the source distribution consists of control cell images, and the target distribution comprises perturbed cell images from the same batch.
@@
To address this distribution-to-distribution problem, CellFlow employs flow matching, a state-of-the-art generative modeling approach designed for distribution-wise transformations (Figure 1d). The framework continuously transforms the source distribution into the target using an ordinary differential equation (ODE) by learning a neural network to approximate a velocity field. This direct and native distribution transformation enabled by flow matching is intuitively more effective than previous methods, which rely on adding extra components to GANs, incorporating the source as a condition, or mapping between distributions and noise using diffusion models (Palma et al., 2023; Hung et al., 2024; Bourou et al., 2024).
@@
We demonstrate the effectiveness of CellFlow on three datasets: BBBC021 (chemical perturbations) (Caie et al., 2010), RxRx1 (genetic modifications via CRISPR or ORF) (Sypetkowski et al., 2023), and JUMP (combined chemical and genetic perturbations) (Chandrasekaran et al., 2023). CellFlow generates high-fidelity images of cellular changes in response to perturbations across all datasets, improving FID scores by 35% over previous approaches. The generated images capture meaningful biological patterns, demonstrated by a 12% improvement in predicting mode-of-action compared to existing methods (Figure 1e). Importantly, CellFlow maintains consistent performance across diverse experimental conditions and generalizes to held-out perturbations never seen during training, showing its broad applicability.
@@
Moreover, CellFlow introduces two key capabilities with significant potential for biological research (Figure 4). First, it effectively corrects batch effects by conditioning on control cells from different batches. By comparing control images with generated images, it can disentangle true perturbation-induced morphological changes from experimental batch artifacts. Second, CellFlow enables bidirectional interpolation between cellular states due to the continuous and reversible nature of the velocity field in flow matching. This interpolation provides a means to explore intermediate cellular morphologies and potentially gain deeper insights into dynamic perturbation responses.
@@
In summary, by formulating cellular morphology prediction as a distribution-to-distribution problem and using flow matching as a solution, CellFlow enables accurate prediction of perturbation responses (Figure 1). CellFlow not only achieves state-of-the-art performance but unlocks new capabilities such as handling batch effects or visualizing cellular state transitions, significantly advancing the field towards a virtual cell for drug discovery and personalized therapy.
@@ 2.2 Data
Figure 2: CellFlow algorithm. (a) Training. The neural network $v_{\theta}$ learns a velocity field by fitting trajectories between control cell images ($x_{0}\sim p_{0}$) and perturbed cell images ($x_{1}\sim p_{1}$). At each training step, intermediate states $x_{t}$ are sampled along the linear interpolation between $x_{0}$ and $x_{1}$, with $t\sim U[0,1]$. The network minimizes the loss $L$, which measures the difference between the predicted velocity $v_{\theta}(x_{t},t,c)$ and the true velocity $(x_{1}-x_{0})$. (b) Inference. The trained velocity field $v_{\theta}$ guides the transformation of a control cell state $x_{0}$ into a perturbed cell state $x_{1}$. This is achieved by solving an ordinary differential equation iteratively, using numerical integration steps over time $t$ (e.g., $t=0.0,0.1,0.2,\ldots,0.9,1.0$). Each step updates the cell state using the learned velocity field.
@@ 3 Method
As detailed in §2, we predict cell morphological changes by transforming distributions between control and perturbed cells under specific conditions within the same batch. In this section, we introduce CellFlow, which leverages flow matching, a principled framework for learning continuous transformations between probability distributions. We adapt flow matching with condition, noise augmentation, and classifier-free guidance to better address our problem setting.
(a) Main Results
BBBC021 (Chemical) RxRx1 (Genetic) JUMP (Combined) Method FIDo FIDc KIDo KIDc FIDo FIDc KIDo KIDc FIDo FIDc KIDo KIDc PhenDiff (MICCAI’24) 49.5 109.2 3.10 3.18 65.9 174.4 5.19 5.29 49.3 127.3 5.09 5.17 IMPA (Nature Comm’25) 33.7 76.5 2.60 2.70 41.6 164.8 2.91 2.94 14.6 99.9 1.08 1.06 CellFlow 18.7 56.8 1.62 1.59 33.0 163.5 2.38 2.40 9.0 84.4 0.63 0.65
(b) Per Perturbation Results
Chemical Perturbations Genetic Perturbations Method Alsterpaullone AZ138 Bryostatin Colchicine Mitomycin C PP-2 ACSS1 CRISP3 RASD1 PhenDiff (MICCAI’24) 106.6 120.0 106.9 111.2 110.0 121.7 157.5 144.6 180.4 IMPA (Nature Comm’25) 69.6 59.9 104.3 84.4 57.0 77.3 152.6 142.7 147.1 CellFlow 41.6 44.4 47.0 72.3 42.3 64.3 140.9 125.1 140.1
Table 1: Evaluation of CellFlow. (a) Main results. CellFlow outperforms GAN- and diffusion-based baselines, achieving state-of-the-art performance in cellular morphology prediction across three chemical, genetic, and combined perturbations datasets. Metrics measure the distance between generated and ground-truth samples, with lower values indicating better performance. FIDo (overall FID) evaluates all images, while FIDc (conditional FID) averages results per perturbation $c$. KID values are scaled by 100 for visualization. (b) Per perturbation results. For six representative chemical perturbations and three genetic perturbations, CellFlow generates significantly more accurate images that better capture the perturbation effects than other methods, as measured by the FID score.
@@ 3.3 Classifier-Free Guidance
We incorporate classifier-free guidance (Ho & Salimans, 2022) to improve generation fidelity. During training, we randomly mask conditions with probability $p_{c}$, replacing $c$ with a null token $\emptyset$. At inference time, we interpolate between conditional and unconditional predictions:
@@ 3.5 Neural Network Architecture
The entire CellFlow algorithm is summarized in §B.
@@
Method MoA Accuracy
Groundtruth Perturbed Image 72.4
PhenDiff 52.6
IMPA 63.7
CellFlow 71.1
(c) Ablation Study
@@
CellFlow w/o Condition 45.0 113.0 2.37 2.37
CellFlow w/o CFG 32.6 92.4 1.23 1.35
CellFlow w/o Noise 31.9 91.4 1.24 1.26
CellFlow 18.7 56.8 1.62 1.59
(b) Out-of-Distribution Generalization
Method FIDo FIDc KIDo KIDc
PhenDiff 67.7 151.6 3.45 3.71
IMPA 44.5 136.9 3.07 3.24
CellFlow 42.0 98.0 1.31 1.23
(d) Batch Effect Study
Method FIDo FIDc KIDo KIDc
CellFlow w/ Other Batch Init 23.7 71.9 2.08 2.09
CellFlow 18.7 56.8 1.62 1.59
Table 2: More evaluation and ablation of CellFlow. (a) MoA classification. On the BBBC021 dataset, we train a classifier to predict the drug’s mode of action (MoA) from cell morphology images and evaluate the accuracy of generated images. CellFlow achieves significantly higher accuracy than other methods, closely aligning with ground-truth images and effectively reflecting the biological effects of perturbations. (b) Out-of-distribution generalization. CellFlow maintains strong performance when generating cell morphology images for novel chemical compounds not seen during training on BBBC021. (c) Ablation study. Removing key components degrades CellFlow’s performance, emphasizing their importance. (d) Batch effect study. CellFlow shows improved performance when using control images from the same batch as initialization, highlighting the critical role of control images in calibrating batch effects.
@@ 4 Results
In this section, we present detailed results demonstrating CellFlow’s state-of-the-art performance in cellular morphology prediction under perturbations, outperforming existing methods across multiple datasets and evaluation metrics.
@@ 4.1 Datasets
Our experiments were conducted using three cell imaging perturbation datasets: BBBC021 (chemical perturbation) (Caie et al., 2010), RxRx1 (genetic perturbation) (Sypetkowski et al., 2023), and the JUMP dataset (combined perturbation) (Chandrasekaran et al., 2023). We followed the preprocessing protocol from IMPA (Palma et al., 2023), which involves correcting illumination, cropping images centered on nuclei to a resolution of 96×96, and filtering out low-quality images. The resulting datasets include 98K, 171K, and 424K images with 3, 5, and 6 channels, respectively, from 26, 1,042, and 747 perturbation types. Examples of these images are provided in Figure 3. Details of datasets are provided in §E.
@@ 4.2 Experimental Setup
Evaluation metrics. We evaluate methods using two types of metrics: (1) FID and KID, which measure image distribution similarity via Fréchet and kernel-based distances, computed on 5K generated images for BBBC021 and 100 randomly selected perturbation classes for RxRx1 and JUMP; we report both overall scores across all samples and conditional scores per perturbation class. (2) Mode of Action (MoA) classification accuracy, which assesses biological fidelity by using a trained classifier to predict a drug’s effect from perturbed images and comparing it to its known MoA from the literature.
Baselines. We compare our approach against two baselines, PhenDiff (Bourou et al., 2024) and IMPA (Palma et al., 2023), the only two baselines that incorporate control images into their model design — a crucial setup for distinguishing true perturbation effects from artifacts such as batch effects. PhenDiff uses diffusion models to first map control images to noise and then transform the noise into target images. In contrast, IMPA employs GANs with an AdaIN layer to transfer the style of control images to target images, specifically designed for paired image-to-image mappings. Our method uses flow matching, which is tailored for distribution-to-distribution mapping, providing a more suitable solution for our problem. We reproduce these baselines with official codes.
Training details. CellFlow employs a UNet-based velocity field with a four-stage design. Perturbations are encoded following IMPA (Palma et al., 2023). Training is conducted for 100 epochs on 4 A100 GPUs. Details are in §C.
Figure 3: Qualitative comparisons. CellFlow generates significantly more accurate images that reflect the actual biological effects of perturbations compared to baselines. For example, Floxuridine inhibits DNA replication, leading to reduced cell density; AZ138 is an Eg5 inhibitor, causing cell death and shrinkage; Demecolcine destabilizes microtubules, resulting in smaller, fragmented nuclei. Columns 1–5, 6–7, and 8–9 correspond to samples from the BBBC021, RxRx1, and JUMP datasets, respectively. More drug’s mode-of-action in §E.

### v3

# CellFlux: Simulating Cellular Morphology Changes via Flow Matching

@@ Abstract
Building a virtual cell capable of accurately simulating cellular behaviors in silico has long been a dream in computational biology. We introduce CellFlux, an image-generative model that simulates cellular morphology changes induced by chemical and genetic perturbations using flow matching. Unlike prior methods, CellFlux models distribution-wise transformations from unperturbed to perturbed cell states, effectively distinguishing actual perturbation effects from experimental artifacts such as batch effects—a major challenge in biological data. Evaluated on chemical (BBBC021), genetic (RxRx1), and combined perturbation (JUMP) datasets, CellFlux generates biologically meaningful cell images that faithfully capture perturbation-specific morphological changes, achieving a 35% improvement in FID scores and a 12% increase in mode-of-action prediction accuracy over existing methods. Additionally, CellFlux enables continuous interpolation between cellular states, providing a potential tool for studying perturbation dynamics. These capabilities mark a significant step toward realizing virtual cell modeling for biomedical research. Project page: https://yuhui-zh15.github.io/CellFlux/.
flow matching, cell image, drug discovery, generative models
@@ 1 Introduction
Figure 1: Overview of CellFlux. (a) Objective. CellFlux aims to predict changes in cell morphology induced by chemical or gene perturbations in silico. In this example, the perturbation effect reduces the nuclear size. (b) Data. The dataset includes images from high-content screening experiments, where chemical or genetic perturbations are applied to target wells, alongside control wells without perturbations. Control wells provide prior information to contrast with target images, enabling the identification of true perturbation effects (e.g., reduced nucleus size) while calibrating non-perturbation artifacts such as batch effects—systematic biases unrelated to the perturbation (e.g., variations in color intensity). (c) Problem formulation. We formulate the task as a distribution-to-distribution problem (many-to-many mapping), where the source distribution consists of control images, and the target distribution contains perturbed images within the same batch. (d) Flow matching. CellFlux employs flow matching, a state-of-the-art generative approach for distribution-to-distribution problems. It learns a neural network to approximate a velocity field, continuously transforming the source distribution into the target by solving an ordinary differential equation (ODE). (e) Results. CellFlux significantly outperforms baselines in image generation quality, achieving lower Fréchet Inception Distance (FID) and higher classification accuracy for mode-of-action (MoA) predictions.
@@
In this work, we introduce CellFlux, an image-generative model that simulates how cellular morphology changes in response to chemical or genetic perturbations (Figure 1a). CellFlux’s key innovation is formulating cellular morphology prediction as a distribution-to-distribution learning problem, and leveraging flow matching (Lipman et al., 2023), a state-of-the-art generative modeling technique designed for distribution-wise transformation, to solve this problem.
@@
Specifically, cell morphology data are collected through high-content microscopy screening, where images of control and perturbed cells are captured from experimental wells across different batches (Figure 1b). Control wells, which receive no drug treatment or genetic modifications, play a crucial role in providing prior information and serving as a reference to distinguish true perturbation effects from other sources of variation. They help calibrate non-perturbation factors, such as batch effects—systematic biases unrelated to perturbations, including variations in color or intensity, akin to distribution shifts in machine learning. Properly incorporating control wells is essential for capturing actual perturbation effects rather than artifacts, yet many existing methods overlook this aspect (Yang et al., 2021; Navidi et al., 2025; Cook et al., 2024). To address this, we frame cellular morphology prediction as a distribution-to-distribution mapping problem (Figure 1c), where the source distribution consists of control cell images, and the target distribution comprises perturbed cell images from the same batch.
@@
To address this distribution-to-distribution problem, CellFlux employs flow matching, a state-of-the-art generative modeling approach designed for distribution-wise transformations (Figure 1d). The framework continuously transforms the source distribution into the target using an ordinary differential equation (ODE) by learning a neural network to approximate a velocity field. This direct and native distribution transformation enabled by flow matching is intuitively more effective than previous methods, which rely on adding extra components to GANs, incorporating the source as a condition, or mapping between distributions and noise using diffusion models (Palma et al., 2025; Hung et al., 2024; Bourou et al., 2024).
@@
We demonstrate the effectiveness of CellFlux on three datasets: BBBC021 (chemical perturbations) (Caie et al., 2010), RxRx1 (genetic modifications via CRISPR or ORF) (Sypetkowski et al., 2023), and JUMP (combined chemical and genetic perturbations) (Chandrasekaran et al., 2023). CellFlux generates high-fidelity images of cellular changes in response to perturbations across all datasets, improving FID scores by 35% over previous approaches. The generated images capture meaningful biological patterns, demonstrated by a 12% improvement in predicting mode-of-action compared to existing methods (Figure 1e). Importantly, CellFlux maintains consistent performance across diverse experimental conditions and generalizes to held-out perturbations never seen during training, showing its broad applicability.
@@
Moreover, CellFlux introduces two key capabilities with significant potential for biological research (Figure 4). First, it effectively corrects batch effects by conditioning on control cells from different batches. By comparing control images with generated images, it can disentangle true perturbation-induced morphological changes from experimental batch artifacts. Second, CellFlux enables bidirectional interpolation between cellular states due to the continuous and reversible nature of the velocity field in flow matching. This interpolation provides a means to explore intermediate cellular morphologies and potentially gain deeper insights into dynamic perturbation responses.
@@
In summary, by formulating cellular morphology prediction as a distribution-to-distribution problem and using flow matching as a solution, CellFlux enables accurate prediction of perturbation responses (Figure 1). CellFlux not only achieves state-of-the-art performance but unlocks new capabilities such as handling batch effects or visualizing cellular state transitions, significantly advancing the field towards a virtual cell for drug discovery and personalized therapy.
@@ 2.2 Data
Figure 2: CellFlux algorithm. (a) Training. The neural network $v_{\theta}$ learns a velocity field by fitting trajectories between control cell images ($x_{0}\sim p_{0}$) and perturbed cell images ($x_{1}\sim p_{1}$). At each training step, intermediate states $x_{t}$ are sampled along the linear interpolation between $x_{0}$ and $x_{1}$, with $t\sim U[0,1]$. The network minimizes the loss $L$, which measures the difference between the predicted velocity $v_{\theta}(x_{t},t,c)$ and the true velocity $(x_{1}-x_{0})$. (b) Inference. The trained velocity field $v_{\theta}$ guides the transformation of a control cell state $x_{0}$ into a perturbed cell state $x_{1}$. This is achieved by solving an ordinary differential equation iteratively, using numerical integration steps over time $t$ (e.g., $t=0.0,0.1,0.2,\ldots,0.9,1.0$). Each step updates the cell state using the learned velocity field.
@@ 3 Method
As detailed in §2, we predict cell morphological changes by transforming distributions between control and perturbed cells under specific conditions within the same batch. In this section, we introduce CellFlux, which leverages flow matching, a principled framework for learning continuous transformations between probability distributions. We adapt flow matching with condition, noise augmentation, and classifier-free guidance to better address our problem setting.
@@ 3.1 Preliminaries
(a) Main Results
BBBC021 (Chemical) RxRx1 (Genetic) JUMP (Combined) Method FIDo FIDc KIDo KIDc FIDo FIDc KIDo KIDc FIDo FIDc KIDo KIDc PhenDiff (MICCAI’24) 49.5 109.2 3.10 3.18 65.9 174.4 5.19 5.29 49.3 127.3 5.09 5.17 IMPA (Nature Comm’25) 33.7 76.5 2.60 2.70 41.6 164.8 2.91 2.94 14.6 99.9 1.08 1.06 CellFlux 18.7 56.8 1.62 1.59 33.0 163.5 2.38 2.40 9.0 84.4 0.63 0.65
(b) Per Perturbation Results
Chemical Perturbations Genetic Perturbations Method Alsterpaullone AZ138 Bryostatin Colchicine Mitomycin C PP-2 ACSS1 CRISP3 RASD1 PhenDiff (MICCAI’24) 106.6 120.0 106.9 111.2 110.0 121.7 157.5 144.6 180.4 IMPA (Nature Comm’25) 69.6 59.9 104.3 84.4 57.0 77.3 152.6 142.7 147.1 CellFlux 41.6 44.4 47.0 72.3 42.3 64.3 140.9 125.1 140.1
Table 1: Evaluation of CellFlux. (a) Main results. CellFlux outperforms GAN- and diffusion-based baselines, achieving state-of-the-art performance in cellular morphology prediction across three chemical, genetic, and combined perturbations datasets. Metrics measure the distance between generated and ground-truth samples, with lower values indicating better performance. FIDo (overall FID) evaluates all images, while FIDc (conditional FID) averages results per perturbation $c$. KID values are scaled by 100 for visualization. (b) Per perturbation results. For six representative chemical perturbations and three genetic perturbations, CellFlux generates significantly more accurate images that better capture the perturbation effects than other methods, as measured by the FID score.
@@ 3.3 Classifier-Free Guidance
We incorporate classifier-free guidance (Ho & Salimans, 2022; Zheng et al., 2023) to improve generation fidelity. During training, we randomly mask conditions with probability $p_{c}$, replacing $c$ with a null token $\emptyset$. At inference time, we interpolate between conditional and unconditional predictions:
@@ 3.5 Neural Network Architecture
The entire CellFlux algorithm is summarized in §B.
@@
Method MoA Accuracy MoA Macro-F1 MoA Weighted-F1
Groundtruth Perturbed Image 72.4 69.7 72.1
PhenDiff 52.6 33.6 52.1
IMPA 63.7 40.2 64.8
CellFlux 71.1 49.0 70.7
(b) Out-of-Distribution Generalization
Method FIDo FIDc KIDo KIDc MoA Accuracy MoA Macro-F1 MoA Weighted-F1
Groundtruth Perturbed Image 0.0 0.0 0.0 0.0 88.0 85.0 88.0
PhenDiff 67.7 151.6 3.45 3.71 9.6 9.3 7.4
IMPA 44.5 136.9 3.07 3.24 16.0 10.0 13.1
CellFlux 42.0 98.0 1.31 1.23 43.2 36.6 42.8
(c) Batch Effect Study
Method FIDo FIDc KIDo KIDc MoA Accuracy MoA Macro-F1 MoA Weighted-F1
CellFlux w/ Other Batch Init 19.9 66.3 1.70 1.69 48.2 32.9 48.4
CellFlux 18.7 56.8 1.62 1.59 71.2 49.0 70.7
(d) Ablation Study
@@
CellFlux w/o Condition 45.0 113.0 2.37 2.37
CellFlux w/o CFG 32.6 92.4 1.23 1.35
CellFlux w/o Noise 31.9 91.4 1.24 1.26
CellFlux 18.7 56.8 1.62 1.59
Table 2: More evaluation and ablation of CellFlux. (a) MoA classification. On the BBBC021 dataset, we train a classifier to predict the drug’s mode of action (MoA) from cell morphology images and evaluate the accuracy/F1 of generated images. CellFlux achieves significantly higher accuracy/F1 than other methods, closely aligning with ground-truth images and effectively reflecting the biological effects of perturbations. (b) Out-of-distribution generalization. CellFlux maintains strong performance when generating cell morphology images for novel chemical compounds not seen during training on BBBC021. (c) Batch effect study. CellFlux shows improved performance when using control images from the same batch as initialization, highlighting the critical role of control images in calibrating batch effects. (d) Ablation study. Removing key components degrades CellFlux’s performance, emphasizing their importance.
@@ 4 Results
In this section, we present detailed results demonstrating CellFlux’s state-of-the-art performance in cellular morphology prediction under perturbations, outperforming existing methods across multiple datasets and evaluation metrics.
@@ 4.1 Datasets
Our experiments were conducted using three cell imaging perturbation datasets: BBBC021 (chemical perturbation) (Caie et al., 2010), RxRx1 (genetic perturbation) (Sypetkowski et al., 2023), and the JUMP dataset (combined perturbation) (Chandrasekaran et al., 2023). We followed the preprocessing protocol from IMPA (Palma et al., 2025), which involves correcting illumination, cropping images centered on nuclei to a resolution of 96×96, and filtering out low-quality images. The resulting datasets include 98K, 171K, and 424K images with 3, 6, and 5 channels, respectively, from 26, 1,042, and 747 perturbation types. Examples of these images are provided in Figure 3. Details of datasets are provided in §E.
@@ 4.2 Experimental Setup
Evaluation metrics. We evaluate methods using two types of metrics: (1) FID and KID (lower the better), which measure image distribution similarity via Fréchet and kernel-based distances, computed on 5K generated images for BBBC021 and 100 randomly selected perturbation classes for RxRx1 and JUMP; we report both overall scores across all samples and conditional scores per perturbation class. (2) Mode of Action (MoA) classification accuracy and F1 score (higher the better), which assesses biological fidelity by using a trained classifier to predict a drug’s effect from perturbed images and comparing it to its known MoA from the literature.
Baselines. We compare our approach against two baselines, PhenDiff (Bourou et al., 2024) and IMPA (Palma et al., 2025), the only two baselines that incorporate control images into their model design — a crucial setup for distinguishing true perturbation effects from artifacts such as batch effects. PhenDiff uses diffusion models to first map control images to noise and then transform the noise into target images. In contrast, IMPA employs GANs with an AdaIN layer to transfer the style of control images to target images, specifically designed for paired image-to-image mappings. Our method uses flow matching, which is tailored for distribution-to-distribution mapping, providing a more suitable solution for our problem. We reproduce these baselines with official codes.
Training details. CellFlux employs a UNet-based velocity field with a four-stage design. Perturbations are encoded following IMPA (Palma et al., 2025). Training is conducted for 100 epochs on 4 A100 GPUs. Details are in §C.
Figure 3: Qualitative comparisons. CellFlux generates significantly more accurate images that reflect the actual biological effects of perturbations compared to baselines. For example, Floxuridine inhibits DNA replication, leading to reduced cell density; AZ138 is an Eg5 inhibitor, causing cell death and shrinkage; Demecolcine destabilizes microtubules, resulting in smaller, fragmented nuclei. Columns 1–5, 6–7, and 8–9 correspond to samples from the BBBC021, RxRx1, and JUMP datasets, respectively. More drug’s mode-of-action in §E.